# Zeyu Shen1    Basileal Imana1    Tong Wu1    Chong Xiang2    Prateek Mittal1    Aleksandra Korolova1

### v1

@@ 5.4 Evaluation Results for Multi-Position Attacks
In this section, we present the evaluation results of both MIS and Sampling + MIS compared against RobustRAG (Keyword) as the baseline on “cleaned” versions of each dataset where we filter out irrelevant documents. We experiment with cleaned datasets because we find that, for the original datasets, most documents retrieved from Google Search as a knowledge-base do not contain the true answer. Figure 2 shows only a small fraction of the documents contain the true answer for questions in the original datasets: For RQA, on average only 26% of the documents provided for a question contain the true answer verbatim, 31% for NQ, and 19% for TQA. Thus, experimenting with cleaned datasets ensures that the study of multi-position attacks is not confounded by pre-existing noise. This allows for a clearer understanding of the direct impact of the number of attacked documents, as any performance degradation can be more confidently attributed to the attack itself rather than the inherent irrelevance of some documents. This also aligns with our theoretical bounds where the number of documents $k$ refers to relevant documents.
@@
Each experiment is repeated 5 times and confidence bands are added. The results are presented in Figure 3 and 4. We observe that MIS and Sampling + MIS typically show a more graceful degradation in performance as more documents are attacked. In contrast, RobustRAG (Keyword) sometimes exhibits sharper drops in accuracy, particularly under a higher number of attacks. For example, for $k=10$, in RQA with 4 attacked documents, MIS maintains an accuracy of around 0.52, while RobustRAG (Keyword) drops to about 0.26. Similar trends are observed for NQ and TQA, where MIS sustains a noticeable advantage as corruption increases. For $k=50$, Sampling + MIS again shows significantly better performance. These results demonstrate the superior robustness of our reliability-aware framework against multi-position suffix attacks.
5.5 Reliability-Awareness of Our Methods
5.6 Running Time Analysis
While any added latency requires justification, it is crucial to consider the context of modern, potentially complex RAG workflows. Simple RAG involves retrieval and a single generation step, but achieving high quality often necessitates more elaborate strategies. Users interacting with sophisticated RAG systems might experience multi-second latencies, which can stem not only from retrieval and basic generation [41, 46] but also from extensive downstream processing, such as reasoning or other test-time scaling techniques applied for enhanced analysis and answer quality.
Our MIS-based approach functions primarily as a document filtering and selection mechanism upstream of this final, potentially costly, answer generation or analysis stage. This contrasts fundamentally with methods such as Keyword Aggregation or Decoding Aggregation [61], which act as alternative inference procedures themselves. The key advantage of our filtering approach is its modularity; it can be seamlessly integrated upstream of any subsequent inference strategy, even though Section 3 presents a specific way that Vanilla RAG is invoked after document selection.
6 Conclusion
Acknowledgments
Appendix A Related Works
Adversarial Attacks Against RAG.
The standard RAG pipeline involves retrieving relevant documents, optionally re-ranking them, and feeding them to the LLM for generation. However, this reliance on external data creates vulnerabilities. Early works studied misinformation attacks against QA models. Recent attacks specifically target LLM-powered RAG, evolving rapidly. Corpus poisoning or retrieval corruption involves injecting malicious content into the knowledge base; examples include PoisonedRAG [70], Topic-FlipRAG [14], MM-PoisonRAG [20]. Prompt injection, embedding malicious instructions in retrieved data, remains a top threat. Other methods include low-level perturbations such as typos (e.g., GARAG [8]) and increasingly, data extraction attacks aiming to steal information from the RAG database using optimization techniques (MARAGE [22]), backdoors implanted during fine-tuning, or automated agent-based methods (RAG-Thief [25]). Some recent works demonstrate the manipulability of LLM preferences concerning products, a significant issue as LLMs are increasingly used for product recommendations. For instance, [29] show that inserting a “strategic text sequence” into product metadata can improve its ranking in LLM outputs, while [45] found that adversarial content on product webpages can alter its own or competitors’ rankings in LLM recommendations.
Robust RAG Frameworks.
Several frameworks aim to improve RAG robustness, addressing resilience against noise or defense against adversarial attacks:
  • RobustRAG: [61] Employs an “isolate-then-aggregate” strategy (using keyword and decoding aggregation) for provable robustness against retrieval corruption. However, their methods have limited scalability to long-form generation. The keyword-based voting, for instance, struggles when answers are lengthy or complex, because it generates the ultimate answer based only on a few keywords and necessarily suffers from significant information loss.
  • InstructRAG: [60] Teaches LLMs to denoise context via self-synthesized rationales, improving robustness to noisy retrievals without extra supervision. However, as shown in our evaluations and [69], it is not robust against simple adversarial attacks such as prompt injection.
  • AstuteRAG: [58] Addresses imperfect retrieval and internal/external knowledge conflicts by eliciting internal knowledge, consolidating sources, and selecting the most reliable answer. It doesn’t explicitly use initial retrieval rank and has been primarily evaluated on short-form QA. Similar to InstructRAG, as shown in our evaluations and [69], it is not robust against simple adversarial attacks such as prompt injection.
  • TrustRAG: [69] Defends against corpus poisoning using $k$-means clustering to filter suspicious documents and LLM self-assessment to resolve conflicts. However, it makes the unrealistic assumption that malicious documents form a separate cluster in the embedding space. This assumption is particularly challenging given that precisely controlling such representations is difficult, and adversaries would naturally strive to craft malicious content to be semantically similar to benign documents to evade detection [70]. In contrast, our assumption centers on the established capabilities of modern NLI models to discern contradictions, even if imperfectly and in adversarial settings, which is a more tenable premise in an adversarial context where attackers prioritize stealth over creating easily detectable patterns.
Document Reweighting, Selection and Filtering.
A complementary line of work seeks to filter or reweigh retrieved passages before any generation occurs, ensuring that the context presented to the language model is already relevant, self-consistent, and trustworthy.
  • Self-RAG: [2] let the LLM emit “reflection” tokens that mark useless documents and trigger additional retrieval. However, it ignores source-level reliability and offers no provable guarantees.
  • Chain-of-Note: [66] has the LLM sequentially read each retrieved document and write a brief “note” assessing its relevance before attempting an answer. If a document is deemed unhelpful in these notes, it can be effectively filtered out and not used in the final answer reasoning.
  • CRAG: [65] introduced a retrieval evaluator that scans the retrieved set and predicts whether the question is answerable with the given documents. If deemed unanswerable, the system can abstain or retrieve from a broader source, rather than force a guess.
  • CrAM: [11] scores each passage with an external credibility estimator and down-weights low-credibility tokens in the LLM’s attention.
  • RA-RAG: [23] Explicitly models source reliability using iterative offline estimation and weighted majority voting aggregation. However, it focuses primarily on reliability estimation instead of aggregation. It has also not been shown to be robust against adversarial attacks.
Sampling-Based Methods for Robustness.
Sampling strategies have long been used to bolster model robustness under adversarial or noisy data conditions. A classic example is Random Sample Consensus (RANSAC) [3], which fits models on randomly sampled data subsets to ignore outliers and find a consensus solution. Modern defenses introduce stochasticity to blunt adversarial attacks. For example, [63] proposes adding random transformations at inference time, which demonstrate effective mitigation of adversarial image perturbations without requiring specialized training.
Recent works have also leveraged sampling for provable robustness. [9] introduces randomized smoothing, which converts any classifier into a certifiably robust one by adding Gaussian noise to inputs and predicting via majority vote. [4] show that a standard reservoir sampling can be made robust to an adversarial input stream by increasing the sample size. Their analysis illustrates how strategic resampling can ensure a representative sample despite an attacker’s attempts to corrupt the data.
Reliability Signals in Ranking and Recommendation Systems.
Reliability signals are also useful in protecting ranking and recommendation systems against spam and manipulation. [19] propagates trust from a small seed of human-vetted pages through the web graph, sharply demoting sites that lie far from trusted regions and reducing the impact of link-spam on search results. [39] shows that embedding explicit user-trust scores into collaborative filtering not only boosts accuracy but also curbs profile-injection attacks, demonstrating the defensive value of trust-weighted aggregation.
Appendix B Theoretical Analysis and Proofs
B.1 Analysis of MIS Robustness with Imperfect NLI
In this section, we present the proof of Theorem 3.1. Note that for the proof, we assume that the errors for each edge occur independently.
Theorem (Theorem 3.1 restated).
Proof B.1 (Proof of Theorem 3.1).
Thus, by union bound, the probability that a malicious document ends up in the maximum independent set is at most $\Pr[\textsc{Bad}_{1}]+\Pr[\textsc{Bad}_{2}]\leq e^{-O(k)}$, so the probability that the maximum independent set does not contain any malicious document is at least $1-e^{-O(k)}$, finishing the proof.
B.1.1 Simulations on Small $k$
(a) $k=10$, $\epsilon_{1}=0.05$, $\epsilon_{2}=0.2$
(b) $k=10$, $\epsilon_{1}=0.05$, $\epsilon_{2}=0.4$
(c) $k=20$, $\epsilon_{1}=0.05$, $\epsilon_{2}=0.2$
(d) $k=20$, $\epsilon_{1}=0.05$, $\epsilon_{2}=0.4$
B.2 Analysis of MIS Robustness with Perfect NLI
Theorem B.2.
Proof B.3.
  1. 1.
  2. 2.
Thus, the set of benign documents $B$ is the maximum independent set.
B.3 Robustness Guarantee for Weighted Sampling
Each of the $T$ rounds of sampling represents an independent Bernoulli trial with success probability $p_{\mathrm{clean}}$, where success means drawing a clean context. Let $C$ be the total number of clean contexts generated across the $T$ rounds. $C$ follows a binomial distribution: $C\sim\operatorname{Binomial}(T,p_{\mathrm{clean}})$. In Theorem B.4, we present the robustness guarantee of Algorithm 2.
Theorem B.4.
$T\geq\frac{1}{2(p_{\mathrm{clean}}-(1-\alpha))^{2}}\log\frac{1}{\delta},$
Proof B.5 (Proof of Theorem B.4).
By Hoeffding’s inequality, we have
Thus, the probability of success is
With $T\geq\frac{1}{2(p_{\mathrm{clean}}-(1-\alpha))^{2}}\log\frac{1}{\delta}$, we have $\Pr[C\geq T(1-\alpha)]\geq 1-\delta$, so Algorithm 2 instantiated with $\mathcal{A}$ is $(1-\delta)\lambda$-robust.
Concrete Instantiation.
Take $\alpha=\frac{1}{2}$, $m=2$ and $\eta=0.1$. Then $p_{\mathrm{clean}}=(1-0.1)^{2}\approx 0.81$. Since $p_{\mathrm{clean}}>1/2$, the condition for the Hoeffding bound is met. With $T=20$ we have a failure probability bound of: $\delta=\exp\left(-2\times 20\times(0.81-0.5)^{2}\right)\approx 0.0214.$ Thus, with these parameters, the algorithm returns a robust answer with probability at least 10.0214=0.97861-0.0214=0.9786, or $97.86\%$.
Appendix C Additional Experimental Setup and Evaluation Results
C.1 Detailed Experimental Setup
Datasets.
We evaluate our methods on both short-answer open-domain question answering (QA) and long-form text generation tasks. For QA, we use RealtimeQA (RQA) [28], Natural Questions (NQ) [32], and TriviaQA (TQA) [27]. For long-form generation, we utilize the Biography generation dataset (Bio) [31]. We use 100 queries from RQA dataset, randomly draw 500 queries from each of NQ dataset and TQA dataset, and 50 queries from Bio dataset. For RQA, we use the 100 queries provided by [61]. This differs from the 500 queries sampled for NQ and TQA because the RobustRAG work, which serves as our primary baseline for comparison, only included this specific set of 100 RQA queries in their Git Repo [62]. This circumvents issues arising from RQA’s real-time nature, as the dataset has not been actively updated recently (latest public data points appear to be from 2023), making it problematic to use current search results for its potentially outdated questions. For each query, we retrieve relevant passages using Google Search. Since we initially retrieve only the search result snippets displayed on the first page, crucial information is often truncated (indicated by ”…”). Consequently, the initial ranking provided by the search engine may not accurately reflect the relevance or quality of the snippet content. To address this, we re-rank the retrieved passages for each query using the mxbai-rerank-large-v2 model [33]. This re-ranking step is a common practice in modern RAG pipelines to enhance context quality and can be performed efficiently.
Evaluation Metrics.
The detailed evaluation metrics we use are:
  • For QA tasks (RQA, NQ, TQA): We assess correctness by comparing the generated answer $r$ against the gold answer $g$. GPT-4o serves as an LLM judge to classify the answer as correct or incorrect based on [68]. The reported metric is Accuracy %, representing the percentage of correctly answered queries.
  • For the long-form Bio generation task: We evaluate the quality of the generated biography $r$ following a multi-aspect LLM-as-a-judge rubric similar to [36]. First, a reference (gold) response $g$ is generated by prompting GPT-4o with the full Wikipedia document of the target person. Subsequently, GPT-4o serves as an LLM judge to compare $r$ against $g$, providing individual scores from 0 to 10 for three distinct criteria: (i) factual accuracy, (ii) relevance and recall, and (iii) coherence and structure. The exact prompt template for grading is presented in Appendix E.2. For each query, these three scores are averaged and scaled to 100. The reported metric, the LLM-Judge Score, is the average of these final per-query scores across all queries evaluated in the dataset.
C.2 Evaluation Results for TQA under Prompt Injection Attack
C.3 Evaluation Results for Corpus Poisoning Attack
MIS ($k=10$).
The MIS method demonstrates strong robustness against poisoning attacks, generally outperforming all other baselines. It mostly maintains better performance, especially on the long-form Bio task, compared to RobustRAG (Keyword). Rank-awareness is evident, with MIS typically performing better when the attack targets Position 10 versus Position 1.
Sampling + MIS ($k=50$).
Similarly, Sampling + MIS shows good resilience. It achieves high empirical accuracy under attacks, particularly when attacks target mid- or lower-ranked documents (positions 25 and 50), reinforcing the effectiveness of reliability-aware sampling framework against poisoning in larger document sets.
C.3.1 Attack Success Rate Versus Attack Position
C.4 Analysis of ReliabilityRAG Parameters
@@
Impact of Varying Context Size ($m$).
With fixed $T=20$ and $\gamma=0.9$, performance generally decreases as $m$ increases when the attack is at Position 1 or Position 25. This is because a larger $m$ increases the likelihood of sampling malicious documents. However, when the attack is at Position 50, where the weight of malicious documents is minimal, performance can improve with a slightly larger $m$. This is because a larger $m$ enables the algorithm to consider more documents, potentially avoiding missing relevant and useful ones.
Impact of Varying Number of Sampling Rounds ($T$).
With fixed $m=2$ and $\gamma=0.9$, performance generally increases with $T$ when the attack is at Position 25 and Position 50. This is substantiated by Theorem B.4, which shows that when $p_{\text{clean}}>1-\frac{k^{\prime}}{k}$, the failure probability decreases exponentially with $T$. In other words, increasing $T$ trades off compute for enhanced robustness. There is little improvement when the attack is at Position 1 though, as the malicious documents carry substantial weight in this scenario (especially after irrelevant documents are filtered out and there can actually be many irrelevant documents among the retrieved ones in our empirical evaluations) and $p_{\text{clean}}$ can be small, so the marginal gains from increasing $T$ are diminished.
Impact of Varying Decay Factor $\gamma$.
With fixed $m=2$ and $T=20$, the choice of $\gamma$ influences the weight distribution across documents. A smaller $\gamma$ concentrates weight on the top-ranked documents and makes the system less robust to attacks targeting higher positions but more resilient to attacks on lower-ranked documents. Conversely, a larger $\gamma$ distributes trust more evenly.
@@
Impact of Varying Weight Decay Scheme.
While our analysis has centered on exponential decay weights ($w(x_{i})\propto\gamma^{i-1}$) — a practical heuristic given that our Google Search retrieved documents lack explicit reliability scores [15] — we also evaluated an alternative linear decay scheme ($w(x_{i})\propto 1-\frac{i}{k}$) for comparison. Figure 9 indicates that linear decay offers slightly enhanced robustness against attack at Position 1, at the cost of marginally reduced robustness for attacks targeting positions 25 and 50. This behavior is a direct consequence of the weight distribution: linear decay assigns a smaller proportion of weight to the highest-ranked documents compared to exponential decay. Both approaches are rank-based heuristics that are reasonable to apply in our small-scale evaluations. In practical deployments, however, the selection of weights should still be informed by an understanding of the reliability landscape, guiding whether to heavily concentrate trust on top-ranked documents or to allocate it more broadly.
C.5 Additional Ablation Studies and Sensitivity Analysis
@@
C.5.1 Evaluation on a Multiple-Choice Dataset
To address potential concerns regarding the use of an LLM-as-a-judge, we conduct experiments on a multiple-choice version of the RealtimeQA dataset. This setup allows for objective, programmatic evaluation. For each question, we created a multiple-choice question with four incorrect options generated by GPT-4o and the one correct ground-truth answer. The results, using Mistral-7B under prompt injection attacks, are presented in Tables 7 and 8. Our methods (MIS and Sampling + MIS) continue to outperform the baselines, demonstrating that their effectiveness is not an artifact of the LLM-based evaluation.
@@ C.5.1 Evaluation on a Multiple-Choice Dataset
C.5.2 Robustness to NLI Degradation
Our theoretical framework accounts for imperfect NLI models, but to empirically test this, we simulate NLI degradation in this section. We repeated the prompt injection attack experiments on RealtimeQA (with Mistral-7B) and artificially inverted the outcome of each NLI contradiction check with a probability $\epsilon$. As shown in Tables 9 and 10, our framework degrades gracefully. Even with $\epsilon=0.5$, our methods remain significantly more robust than Vanilla RAG, demonstrating that the defense does not catastrophically fail even when the NLI signal is heavily corrupted.
@@ C.5.2 Robustness to NLI Degradation
C.5.3 Sensitivity to NLI Model and Contradiction Threshold $\beta$
@@
C.6 Evaluation on Adaptive Attack
@@
To verify the intuition that the adaptive attack we consider, though more likely to slip through the contradiction checking and MIS-based filtering, is less harmful, we test the performance of VanillaRAG under the adaptive and non-adaptive attack for each attack position, respectively. The results, as presented in Table 14, show that the adaptive attack is indeed not as harmful as the non-adaptive attack. This phenomenon echoes the “jailbreak tax” identified by [47], which shows that guardrail–bypassing prompts typically suffer a marked drop in downstream utility. Hence, although the adaptive attack helps the malicious “A or B” document slip into the MIS more often, its reduced utility means overall answer accuracy remains largely unchanged.
@@ D.1.1 Discussion of Weight Selection
An intuitive choice is exponentially decaying weights, where $w(x_{i})\propto\gamma^{i-1}$ for some decay factor $0<\gamma<1$, normalized so that $\sum_{i=1}^{k}w(x_{i})=1$. This scheme assigns significantly more importance to top-ranked documents. Such exponential weighting is frequently employed in time series analysis (Exponentially Weighted Moving Average [37]) to give more influence to recent data points, analogous to giving more influence to higher-ranked documents. While sometimes adopted for simplicity and its practical fit to data rather than strict theoretical derivation in some domains, exponential weighting is a well-established technique for incorporating recency or priority into aggregate measures. Choosing an appropriate $\gamma$ often involves balancing the desire to emphasize top documents against the need to retain information from lower-ranked ones.

### v2

@@ C.5 Full Evaluation Results for Multi-Position Attacks
In this section, we present the full evaluation results of both MIS and Sampling + MIS compared against RobustRAG (Keyword) as the baseline on “cleaned” versions of each dataset where we filter out irrelevant documents. We experiment with cleaned datasets because we find that, for the original datasets, most documents retrieved from Google Search as a knowledge-base do not contain the true answer. Figure 6 shows only a small fraction of the documents contain the true answer for questions in the original datasets: For RQA, on average only 26% of the documents provided for a question contain the true answer verbatim, 31% for NQ, and 19% for TQA. Thus, experimenting with cleaned datasets ensures that the study of multi-position attacks is not confounded by pre-existing noise. This allows for a clearer understanding of the direct impact of the number of attacked documents, as any performance degradation can be more confidently attributed to the attack itself rather than the inherent irrelevance of some documents. This also aligns with our theoretical bounds where the number of documents $k$ refers to relevant documents.
@@
Each experiment is repeated 5 times and confidence bands are added. The results are presented in Figure 7 and 8. We observe that MIS and Sampling + MIS typically show a more graceful degradation in performance as more documents are attacked. In contrast, RobustRAG (Keyword) sometimes exhibits sharper drops in accuracy, particularly under a higher number of attacks. For example, for $k=10$, in RQA with 4 attacked documents, MIS maintains an accuracy of around 0.52, while RobustRAG (Keyword) drops to about 0.26. Similar trends are observed for NQ and TQA, where MIS sustains a noticeable advantage as corruption increases. For $k=50$, Sampling + MIS again shows significantly better performance. These results demonstrate the superior robustness of our reliability-aware framework against multi-position suffix attacks.
C.6 Analysis of ReliabilityRAG Parameters
@@
Impact of Varying Context Size ($m$). With fixed $T=20$ and $\gamma=0.9$, performance generally decreases as $m$ increases when the attack is at Position 1 or Position 25. This is because a larger $m$ increases the likelihood of sampling malicious documents. However, when the attack is at Position 50, where the weight of malicious documents is minimal, performance can improve with a slightly larger $m$. This is because a larger $m$ enables the algorithm to consider more documents, potentially avoiding missing relevant and useful ones.
Impact of Varying Number of Sampling Rounds ($T$). With fixed $m=2$ and $\gamma=0.9$, performance generally increases with $T$ when the attack is at Position 25 and Position 50. This is substantiated by Theorem 3, which shows that when $p_{\text{clean}}>1-\frac{k^{\prime}}{k}$, the failure probability decreases exponentially with $T$. In other words, increasing $T$ trades off compute for enhanced robustness. There is little improvement when the attack is at Position 1 though, as the malicious documents carry substantial weight in this scenario (especially after irrelevant documents are filtered out and there can actually be many irrelevant documents among the retrieved ones in our empirical evaluations) and $p_{\text{clean}}$ can be small, so the marginal gains from increasing $T$ are diminished.
Impact of Varying Decay Factor $\gamma$. With fixed $m=2$ and $T=20$, the choice of $\gamma$ influences the weight distribution across documents. A smaller $\gamma$ concentrates weight on the top-ranked documents and makes the system less robust to attacks targeting higher positions but more resilient to attacks on lower-ranked documents. Conversely, a larger $\gamma$ distributes trust more evenly.
Impact of Varying Weight Decay Scheme. While our analysis has centered on exponential decay weights ($w(x_{i})\propto\gamma^{i-1}$) — a practical heuristic given that our Google Search retrieved documents lack explicit reliability scores Google (2024) — we also evaluated an alternative linear decay scheme ($w(x_{i})\propto 1-\frac{i}{k}$) for comparison. Figure 10 indicates that linear decay offers slightly enhanced robustness against attack at Position 1, at the cost of marginally reduced robustness for attacks targeting positions 25 and 50. This behavior is a direct consequence of the weight distribution: linear decay assigns a smaller proportion of weight to the highest-ranked documents compared to exponential decay. Both approaches are rank-based heuristics that are reasonable to apply in our small-scale evaluations. In practical deployments, however, the selection of weights should still be informed by an understanding of the reliability landscape, guiding whether to heavily concentrate trust on top-ranked documents or to allocate it more broadly.
@@
C.7 Additional Ablation Studies and Sensitivity Analysis
@@
C.7.1 Evaluation on a Multiple-Choice Dataset
To address potential concerns regarding the use of an LLM-as-a-judge, we conduct experiments on a multiple-choice version of the RealtimeQA dataset. This setup allows for objective, programmatic evaluation. For each question, we created a multiple-choice question with four incorrect options generated by GPT-4o and the one correct ground-truth answer. The results, using Mistral-7B under prompt injection attacks, are presented in Tables 6 and 7. Our methods (MIS and Sampling + MIS) continue to outperform the baselines, demonstrating that their effectiveness is not an artifact of the LLM-based evaluation.
@@ C.7.1 Evaluation on a Multiple-Choice Dataset
C.7.2 Robustness to NLI Degradation
Our theoretical framework accounts for imperfect NLI models, but to empirically test this, we simulate NLI degradation in this section. We repeated the prompt injection attack experiments on RealtimeQA (with Mistral-7B) and artificially inverted the outcome of each NLI contradiction check with a probability $\epsilon$. As shown in Tables 8 and 9, our framework degrades gracefully. Even with $\epsilon=0.5$, our methods remain significantly more robust than Vanilla RAG, demonstrating that the defense does not catastrophically fail even when the NLI signal is heavily corrupted.
@@ C.7.2 Robustness to NLI Degradation
C.7.3 Sensitivity to NLI Model and Contradiction Threshold $\beta$
@@
C.8 Evaluation on Adaptive Attack
@@
To verify the intuition that the adaptive attack we consider, though more likely to slip through the contradiction checking and MIS-based filtering, is less harmful, we test the performance of VanillaRAG under the adaptive and non-adaptive attack for each attack position, respectively. The results, as presented in Table 13, show that the adaptive attack is indeed not as harmful as the non-adaptive attack. This phenomenon echoes the “jailbreak tax” identified by nikolić2025jailbreaktaxusefuljailbreak, which shows that guardrail–bypassing prompts typically suffer a marked drop in downstream utility. Hence, although the adaptive attack helps the malicious “A or B” document slip into the MIS more often, its reduced utility means overall answer accuracy remains largely unchanged.
@@ D.1.1 Discussion of Weight Selection
An intuitive choice is exponentially decaying weights, where $w(x_{i})\propto\gamma^{i-1}$ for some decay factor $0<\gamma<1$, normalized so that $\sum_{i=1}^{k}w(x_{i})=1$. This scheme assigns significantly more importance to top-ranked documents. Such exponential weighting is frequently employed in time series analysis (Exponentially Weighted Moving Average Luxenberg and Boyd (2024)) to give more influence to recent data points, analogous to giving more influence to higher-ranked documents. While sometimes adopted for simplicity and its practical fit to data rather than strict theoretical derivation in some domains, exponential weighting is a well-established technique for incorporating recency or priority into aggregate measures. Choosing an appropriate $\gamma$ often involves balancing the desire to emphasize top documents against the need to retain information from lower-ranked ones.